OpenStudy (anonymous):

Solve by factoring 5x^3 - 80x=0

jagr2713 (jagr2713):

is there an anser choice

OpenStudy (anonymous):

no

jagr2713 (jagr2713):

i got this x=4

OpenStudy (anonymous):

oh wait my teacher got -4,0,4

OpenStudy (matt101):

\[5x^3-80x=0\]\[5x(x^2-16)=0\]\[5x(x-4)(x+4)=0\]\[x=0,4,-4\]

jagr2713 (jagr2713):

same thing i got

OpenStudy (isaiah.feynman):

@jagr2713 This is a cubic equation, therefore you should have three solutions not just one.

jagr2713 (jagr2713):

oh i got this look

jagr2713 (jagr2713):

Factor out the GCF of 5x from the expression 5x3. 5x(x2)−80x=0 Factor out the GCF of 5x from the expression −80x. 5x(x2)+5x(−16)=0 Factor out the GCF of 5x from 5x3−80x. 5x(x2−16)=0 The binomial can be factored using the difference of squares formula, because both terms are perfect squares. 5x(x+4)(x−4)=0 5x(x+4)(x−4)=0 Set the single term factor on the left-hand side of the equation equal to 0. 5x=0 Divide each term in the equation by 5. More Detail Divide each term in the equation by 5. 5x5=05 Cancel the common factor of 5 in 5x5. 1 5 x 5 =05 Reduce the expression 5x5 by removing a factor of 5 from the numerator and denominator. x=05 0 divided by any number or variable is 0. x=0 x=0 Set each of the factors of the left-hand side of the equation equal to 0. x+4=0 x−4=0 Solve the equations for x. Since 4 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 4 from both sides. x=−4 x−4=0 Set each of the factors of the left-hand side of the equation equal to 0. x=−4 x−4=0 Since −4 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4 to both sides. x=−4 x=4 Solve the equations for x. x=0,−4,4 x=0,−4,4